# Indoor Location Fingerprinting Privacy: A Comprehensive Survey

### v1

@@ 5.2.5. Probability-based attack
5.2.6. Context linking attack
@@
Most location-based attacks make use of some context. Correlating metadata from different sources, such as social media profiles, public databases, or online activities, allows attackers to link contexts and uncover additional details about individuals. When launching a localization attack, it is simple to combine contextual knowledge with the observed location data to determine a target’s exact location. For instance, a personal context linking attack (Gruteser and Grunwald, 2003) may be used to remove all the unnecessary areas from a user’s location after reducing an obfuscated area to a specified spot. Contextual information can be used in conjunction with precise location data to carry out identity attacks. An attacker can assume that their target is or was in a special room of a hospital at a particular moment, for instance, if they know someone’s affiliation and discover it on a hospital room list.
5.2.7. Data linking attack
@@
It is introduced in (Fathalizadeh et al., 2022) in whcih attacks on published indoor location datasets are considered twofold: identification/re-identification and Record/Attribute/Table linkage attacks. The attacker’s goal in an identification/re-identification attack is to identify/re-identify the victim in the whole dataset among all the other users that might be presented there. In an indoor location data set, an attacker with some background knowledge about an individual can identify whether he/she has been in a specific location. This is particularly important when spatial-temporal locations and trajectories are monitored by an adversary. A linkage attack occurs when an adversary can link a record owner (i) to a record in a published dataset, (ii) to a sensitive attribute in a revealed dataset, or (iii) to the published table of data by itself. These attacks are called record linkage, attribute linkage, and table linkage attacks, respectively. Linkage attacks are characterized by the adversary’s prior knowledge. In the case of record linkage and attribute linkage, the adversary knows that a specific individual’s data is presented in a dataset and he wants to learn about sensitive information. Table linkage attack, on the other hand, focuses on understanding whether a known individual’s information is available in the released dataset or not. The adversary could deduce sensitive information from the disclosed dataset based on the distribution of sensitive values in the group to which the individual belongs in an attribute linkage attack
6. Indoor Location Fingerprinting Privacy Preservation Mechanisms (ILFPPM)
@@
This section reviews existing studies on ILFPPM. Four major privacy preservation techniques have been utilized for this purpose, including cryptographic methods, anonymization techniques, Differential Privacy (DP), and Federated Learning (FL). Each of these methods is discussed in the following sub-sections, accompanied by a brief explanation of the privacy method, necessary definitions, applicability of privacy methods to indoor settings, as well as their respective advantages and disadvantages.
@@
Note that there are also passive solutions for localization (Schauer et al., 2016; Gschwandtner and Schindhelm, 2011), in which the fingerprinting dataset is available for the users and the localization process is done locally on the user’s device. In this method, however, the privacy of the user is preserved, but the server-side privacy, or in other words, server data privacy is not preserved, so it is not considered in this categorization.
6.1. Cryptographic Methods
@@
Here, we briefly explain the cryptographic methods and then discuss the studies in which these methods are used in the concept of ILFPPM.
6.1.1. Cryptographic techniques
Definition 6.1 (Homomorphic encryption).
@@ 6.1.1. Cryptographic techniques
Homomorphic encryption is a cryptographic technique that enables computations to be executed on encrypted data directly, without the need for prior decryption. The outcomes of these computations remain encrypted, and upon decryption, yield results identical to those obtained had the operations been conducted on the original, unencrypted data. Two well-known homomorphic-based cryptosystems used in protecting the privacy of indoor location data are DGK (Damgard et al., 2008) and Paillier (Paillier, 1999). The DGK protocol was created by Damgård, Geisler, and Krøigaard in 2007 as an efficient solution to solving the millionaire’s problem. This protocol is a homomorphic cryptosystem dedicated to small plaintexts and fits nicely within the secure comparison protocol. The Paillier cryptosystem is a probabilistic asymmetric algorithm for public key cryptography. The scheme is an additive homomorphic cryptosystem; this means that, given only the public key and the encryption of $m_{1}$ and $m_{2}$, one can compute the encryption of $m_{1}+m_{2}$. It also allows the multiplication of a ciphertext by a plaintext number.
Figure 8. Homomorphic encryption vs SMPC in indoor location fingerprinting systems.
Definition 6.2 (Multi-Party Computations (MPC)).
Secure Multi-Party Computation (SMPC), or simply Multi-Party Computation (MPC), is a subfield within cryptography that aims to develop techniques enabling multiple parties to collaboratively compute a function using their inputs while maintaining the privacy of those inputs.
@@
MPC protocols, especially in scenarios with unconditional private channels, rely on secret sharing. In secret-sharing-based (or Aritmethic secret sharing) methods (Shamir, 1979), the participating parties do not have specific roles, as seen in the creator-evaluator paradigm of Yao’s protocol. Instead, the data associated with each wire is shared among the parties, and a protocol is then employed to evaluate each gate. In this context, the function is defined as a ‘circuit’ over a finite field, unlike the binary circuits used in Yao’s protocol. Such circuits are referred to as ‘arithmetic circuits’ in the literature and consist of addition and multiplication ‘gates’ where the values being operated on are defined over a finite field.
Table 1. Studies on cryptography-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model (Li et al., 2014) 2014 kNN Paillier Simulated Private Coms. Overhead Loc. Error Fully Malicious Loc. & Data Priv. Attack (Wang et al., 2015) 2015 Fuzzy Logic Paillier CRAWDAD Loc. Error Semi-trusted Loc. Priv. Attack (Zhang et al., 2016) 2016 SVM Paillier Private Coms. Overhead Fully malicious Loc. & Data Priv. Attack (Yang and Järvinen, 2018) 2018 - Paillier Garbled Circuit - - Semi-trusted Loc. & Data Priv. Attack Protocol-based Attack (Yang and Järvinen, 2018) 2018 - Paillier Garbled Circuit Simulated - Unilateral-malicious Loc. & Data Priv. Attack Protocol-based Attack (Richter et al., 2018) 2018 kNN Paillier Garbled Circuit Private JUIndoorLoc Tampere Uni Loc. Error Floor Detection - Loc. Priv. Attack (Quijano and Akkaya, 2019) 2019 kNN Paillier DGK Algorithm Colombia Uni Execution Time Battery Life - - (Wang et al., 2019) 2019 kNN ABY Framework Arithmetic Sharing Boolean Sharing Yao’s Sharing Simulated Private same as (Li et al., 2014) Coms. Overhead Run Time Fully malicious Protocol-based attack (Järvinen et al., 2019) 2019 kNN Trusted Third Party Simulated Execution Time Energy Consumption Fully trusted Loc. & Data Priv. Attack Protocol-based Attack (Eshun and Palmieri, 2019) 2019 kNN Paillier Spatial Bloom - Coms. Overhead Semi-trusted Loc. Priv. Attack (Zhang et al., 2020) 2020 kNN Paillier Private Coms. Overhead Loc. Error Time Cost Fully malicious Loc. & Data Priv. Attack (Wu et al., 2020) 2020 kNN Expectation- maximization Bayes Network UJIIndoorLoc Coms. Overhead Loc. Error Time Cost Semi-trusted Loc. Priv. Attack (Nieminen and Järvinen, 2021) 2021 kNN Paillier One Time Pad Garbled Circuits Simulated Uni of Helsinki Coms. Overhead Loc. Error Time Cost Semi-trusted Loc. & Data Priv. Attack (van der Beets et al., 2022) 2022 kNN Arithmetic Sharing Delta Sharing Yao’s Sharing Simulated Coms. Overhead Run Time Semi-trusted - (Hu et al., 2022) 2022 MLE Paillier Private Coms. Overhead Loc. Error - - (Wang et al., 2024) 2024 kNN Inner Product Encryption Simulated JUIndoorLoc Loc. Error Coms. Overhead Time Cost Semi-trusted Data Priv. Attack
6.1.2. Cryptography-based ILFPPM

### v2

@@ 4.2.5 Probability-based attack
4.2.6 Context linking attack
@@
Most location-based attacks make use of some context. Correlating metadata from different sources, such as social media profiles, public databases, or online activities, allows attackers to link contexts and uncover additional details about individuals. When launching a localization attack, it is simple to combine contextual knowledge with the observed location data to determine a target’s exact location. For instance, a personal context linking attack [50] may be used to remove all the unnecessary areas from a user’s location after reducing an obfuscated area to a specified spot. Contextual information can be used in conjunction with precise location data to carry out identity attacks. An attacker can assume that their target is or was in a special room of a hospital at a particular moment, for instance, if they know someone’s affiliation and discover it on a hospital room list.
4.2.7 Data linking attack
@@
Attacks on the published indoor location datasets are considered  [49]: identification/re-identification and Record/Attribute/Table linkage attacks. The attacker’s goal in the identification/re-identification attack is to identify or re-identify the victim in the whole dataset among all the other users that might be presented there. In an indoor location data set, an attacker with some background knowledge about an individual can identify whether he/she has been in a specific location. This is particularly important when spatial-temporal locations and trajectories are monitored by an adversary. A linkage attack occurs when an adversary can link a record owner (i) to a record in a published dataset, (ii) to a sensitive attribute in a revealed dataset, or (iii) to the published table of data by itself. These attacks are called record linkage, attribute linkage, and table linkage attacks, respectively. Linkage attacks are characterized by the adversary’s prior knowledge. In the case of record linkage and attribute linkage, the adversary knows that a specific individual’s data is presented in a dataset and he wants to learn about sensitive information. Table linkage attack, on the other hand, focuses on understanding whether a known individual’s information is available in the released dataset or not. The adversary could deduce sensitive information from the disclosed dataset based on the distribution of sensitive values in the group to which the individual belongs in an attribute linkage attack
5 Indoor Location Fingerprinting Privacy Preservation Mechanisms (ILFPPM)
@@
This section reviews and compares existing studies on ILFPPM. We first discuss the evaluation metrics and the datasets considered in these studies. Four major privacy preservation techniques have been utilized for ILFPPM, including cryptographic methods, anonymization techniques, DP, and FL. Each of these methods is discussed in the following subsections, accompanied by a brief explanation of the privacy method, necessary definitions, the applicability of privacy methods to indoor settings, as well as their respective advantages and disadvantages.
@@
Note that there are also passive solutions for localization [51, 52], in which the fingerprinting dataset is available for the users and the localization process is done locally on the user’s device. In this method, however, the privacy of the user is preserved, but the server-side privacy, or in other words, server data privacy is not preserved, so it is not considered in this categorization.
5.1 Evaluation Metrics
Here, we discuss various metrics used in the different studies to evaluate the ILFPPM. All these metrics and their associated research are mentioned in Tables 3, 4, 5, 6. We categorize all metrics into the three main groups: localization accuracy, Quality of Service (QoS), and privacy metrics.
5.1.1 Localization accuracy metrics
The widely used metric in ILFPPM determines the localization accuracy. Location error metrics are crucial for evaluating the accuracy and precision of localization systems in various applications. These metrics are referred to by the following names: Location Error, Distance Error, Root Mean Square Error (RMSE), Mean Squared Error (MSE), Mean Absolute Error (MAE), Accuracy, Error Rate, Success Rate, Spatial Loss, RSS Loss, Cloaked Region (CR) Area, Area under the Success Rate curve (Area of ASR), Floor Detection, Kendall’s Tau Distance, Control Value, Count, and Area.
1) Location Error and Distance Error are basic metrics that measure the Euclidean distance between the estimated location and the actual location. These metrics help assess how far off the localization predictions are in real-world scenarios.
2) Root Mean Square Error (RMSE) is the square root of the average of the squared differences between the predicted and actual locations. RMSE provides a comprehensive measure of the localization error by giving more weight to larger errors, making it useful for applications where large deviations are especially problematic.
3) Mean Squared Error (MSE) is the average of the squared differences between predicted and actual locations. It penalizes larger errors more than smaller ones, making it effective in scenarios where minimizing larger errors is a priority.
4) Mean Absolute Error (MAE) is the average of the absolute differences between the predicted and actual locations. It provides a straightforward measure of the average magnitude of localization errors, without penalizing larger errors more heavily than smaller ones, as MSE does.
5) Accuracy, Error Rate, and Success Rate are metrics that represent the percentage of correctly localized instances. Accuracy measures the proportion of correct predictions, Error Rate quantifies the percentage of incorrect localizations, and Success Rate refers to the ratio of successful localization attempts (those within a certain error threshold) to the total number of attempts, offering a simple way to evaluate system performance.
6) Spatial Loss measures the discrepancy between the predicted and actual locations in a spatial context, considering the geometric relationships between points. This metric is particularly important when dealing with spatial data, where the relative positions of locations are critical.
7) RSS Loss evaluates the difference in the Received Signal Strength (RSS) values between predicted and actual locations. This metric is often used in localization systems that rely on wireless signal strength measurements to infer location.
8) Cloaked Region (CR) Area refers to the size of the area in which a user’s location is obfuscated to preserve privacy. A larger CR area implies greater uncertainty about the user’s precise location, which can enhance privacy but potentially reduce localization accuracy.
9) Area of ASR (Area under the Success Rate curve) represents the total area under the curve that plots the success rate over varying error thresholds. It provides an aggregate measure of how well the system performs over a range of localization error tolerances.
10) Floor Detection Accuracy is specific to indoor localization systems that operate across multiple floors of a building. It measures the system’s ability to correctly identify the floor on which the user is located.
11) Kendall’s Tau Distance is a ranking metric used to evaluate the correlation between the predicted and actual orderings of locations. It is relevant in systems that rank possible locations or trajectories.
12) Control Value refers to a metric that adjusts or controls for certain variables during localization, ensuring that the system’s performance is assessed under standardized conditions.
13) Count and Area metrics are more context-specific, with Count typically referring to the number of successful localization attempts or the number of messages exchanged, and Area referring to spatial areas of interest, such as the size of the localization region or the area affected by localization errors.
5.1.2 Quality of Service (QoS) metrics
The second group is employed to assess the Quality of Service (QoS) of the ILFPPM. It includes various metrics, such as QoS Loss, Communication Overhead, Time Cost, Run Time, Execution Time, Response Time, Energy Consumption, Entropy, Number of Messages, Noise Effect, Bandwidth Cost, and Battery Life.
1) QoS Loss measures the degradation in service quality concerning predetermined thresholds, indicating the system’s ability to meet desired performance levels. A higher QoS loss indicates that the system is unable to maintain the required performance, especially under varying loads or adverse conditions.
2) Computation and Communication Overhead quantify the additional computational or communication resources required to execute tasks. This is particularly relevant in cryptography-based approaches, where encryption introduces extra processing requirements and data transmission overheads. These overheads, while protecting data privacy and security, increase resource consumption, potentially reducing efficiency and responsiveness.
3) Time Cost metrics encompass various aspects, such as Execution Time, which refers to the duration it takes to complete a specific task, and Run Time, which measures the total duration of system operation. These collectively shape the system’s responsiveness and real-time performance, indicating how well the system can function under different conditions or workloads.
4) Response Time refers to the time it takes for the system to respond to a query or input. In real-time systems, low response time is crucial for maintaining user satisfaction and ensuring the system is reactive to environmental changes.
5) Energy Consumption measures the amount of energy required to execute a task or keep the system running. This is a critical metric, especially for battery-powered devices, as excessive energy consumption can reduce device longevity and make the system impractical for mobile or IoT devices in continuous operation.
6) Entropy is a measure of uncertainty or randomness in the system. In the context of ILFPPM, it can refer to the degree of uncertainty in localization data or how well the system maintains data privacy by introducing uncertainty into location predictions to prevent inference attacks.
7) Number of Messages exchanged during communication sessions provides insights into network traffic and resource utilization. A high number of messages can increase communication overhead, impacting both system scalability and efficiency, particularly in environments with limited bandwidth or high latency.
8) Noise Effect metrics evaluate the impact of signal distortion or interference on data transmission and processing accuracy. Noise can come from various sources, such as physical obstacles, environmental interference, or network congestion. Assessing the noise effect is essential for ensuring robustness and system reliability, especially in challenging, noisy environments.
9) Bandwidth Cost refers to the amount of bandwidth consumed during data transmission. Systems that require frequent communication or transfer large amounts of data can impose significant bandwidth costs, which can affect both the user experience and network efficiency, especially in low-bandwidth scenarios.
10) Battery Life evaluates the longevity of devices operating in the system. Systems that consume excessive battery power may limit the practicality of ILFPPM, particularly for mobile or IoT devices, where energy efficiency is a key requirement.
11) Adaptation speed measures how quickly a model adjusts to new environments by either measuring the number of training steps needed to reach a target accuracy or assessing accuracy after a fixed number of steps. It is calculated as the inverse of these values, offering a way to evaluate how efficiently a model adapts to changes.
5.1.3 Privacy metrics
The last group consists of privacy-related metrics. Privacy metrics are essential for evaluating the effectiveness of privacy-preserving mechanisms in data handling, storage, and transmission. These metrics are Location Privacy, Probabilistic Metrics, and Normalized Central Penalty (NCP) in ILFPPM papers.
1) Location Privacy Metrics evaluate the degree to which a system preserves the location privacy of individuals when handling spatial datasets. These metrics often consider aspects like spatial granularity, where the system determines the precision of location data shared, and proximity disclosure, which measures the risk of revealing sensitive locations based on proximity to known points of interest. Higher granularity or more frequent proximity disclosures can lead to increased privacy risks, making these metrics vital for understanding the trade-offs between utility and privacy in localization systems. Techniques such as $k$-anonymity, location obfuscation, or cloaking are often evaluated using these metrics to determine their effectiveness in safeguarding location privacy.
2) Probabilistic Metrics are particularly important for privacy-preserving mechanisms based on DP. These metrics assess the likelihood that an adversary can infer sensitive information from seemingly anonymized or noisy data. In the context of DP-based indoor localization systems, probabilistic metrics measure the probability of identifying an individual or location through repeated queries or by analyzing the noise added to the location data. By quantifying the risks associated with probabilistic inference attacks, these metrics help in determining whether a system adheres to the privacy guarantees promised by differential privacy. They also highlight the trade-off between data utility (how useful the data is) and privacy protection, ensuring the system effectively balances the two.
3) Normalized Central Penalty (NCP) is a quantitative measure used to assess privacy violations in datasets, often applied to anonymized or perturbed data. It indicates how much sensitive information is leaked in a dataset by evaluating the centrality or proximity of sensitive records. Higher NCP values indicate greater privacy risks, showing that more sensitive data is potentially exposed. In the context of ILFPPM, NCP is used to analyze the extent to which individual locations or other sensitive attributes can be reidentified after anonymization or perturbation. By providing a numerical measure of privacy leakage, NCP allows researchers and developers to evaluate the effectiveness of privacy-preserving mechanisms and compare them across different systems. This is especially useful when determining the trade-off between privacy protection and system performance, helping to fine-tune privacy mechanisms to achieve optimal results.
Other important privacy metrics include Adversary Success Rate, which measures the likelihood that an attacker can successfully breach the system’s privacy defenses. A lower adversary success rate corresponds to a more secure system. Anonymity Set Size measures how many individuals share the same privacy profile, with larger sets providing better anonymity guarantees. Exposure Risk evaluates the likelihood that sensitive information is exposed over time through repeated queries or location data sharing. Finally, Linkability measures the ability of an adversary to link anonymized records back to the original individual or location, thus breaching privacy. High linkability scores indicate a vulnerability in the anonymization or obfuscation techniques used by the system.
5.2 Indoor Location Datasets
Given the diversity of proposed methods and the fact that many of them may not be highly applicable to real-world datasets, most papers, even those intended for application to real-world datasets, employ their proposed methods in a simulated testbed. The open-source datasets widely used for indoor localization, even in non-privacy-focused papers, are UJIndoorLoc [53] and JUIndoorLoc [54]. The CSUCY [55], KIOSUCI [55], CRAWDAD [56], PosData [57], SPAWC [58], Geolife, and Gowalla [41] are also public datasets that are used in ILFPPM papers. In some papers, the dataset’s location is not explicitly mentioned; instead, they mention the physical properties of the real-world dataset prepared for the proposed method. Some other datasets are Cyberspace Research Institute (CRI) [59], Sangmyung University [60], Rutgers University [61], Tampere University [62], Colombia University [63], University of Helsinki [64], University of Minho [65], and Guangzhou Xinguang shopping mall [66] datasets.
5.3 Cryptography in ILFPPM
@@
Here, we briefly explain the cryptographic methods and then discuss the studies in which these methods are used in the concept of ILFPPM. Finally, we provide guidelines on utilizing cryptography in ILFPPM.
5.3.1 Cryptographic techniques
@@
Homomorphic encryption (HE) Homomorphic encryption is a cryptographic technique that enables computations to be executed on encrypted data directly, without the need for prior decryption. The outcomes of these computations remain encrypted, and upon decryption, yield results identical to those obtained had the operations been conducted on the original, unencrypted data. Two well-known homomorphic cryptosystems used to protect the privacy of indoor location data are DGK [67] and Paillier [68]. The DGK protocol, developed by Damgård, Geisler, and Krøigaard in 2007, is an efficient solution for the millionaire’s problem and supports homomorphic operations on small plaintexts. Similar to Paillier, DGK allows computations on encrypted data, but the plaintext space in DGK is smaller and can be chosen dynamically. The computations in DGK are performed modulo $N$, whereas in Paillier they are performed modulo $N^{2}$. The Paillier cryptosystem is an additive homomorphic cryptosystem, meaning that given the encryption of two values, $m_{1}$ and $m_{2}$, one can compute the encryption of their sum, $m_{1}+m_{2}$. Additionally, Paillier supports the multiplication of a ciphertext by a plaintext number.
Figure 8: Homomorphic encryption vs SMPC in ILF systems.
@@
Secure Multi-Party Computation (SMPC): Secure Multi-Party Computation (SMPC), or simply Multi-Party Computation (MPC), is a subfield within cryptography that aims to develop techniques enabling multiple parties to collaboratively compute a function using their inputs while maintaining the privacy of those inputs. MPC protocols can rely on either secret sharing or garbled circuits. In secret-sharing-based methods [69], the participating parties do not have specific roles, unlike the garbler-evaluator paradigm seen in Yao’s protocol. Instead, the data associated with each wire is shared among the parties, and a protocol is employed to evaluate each gate. The function is defined as a ‘circuit’ over a finite field, although it can also be defined over a ring or bits, as seen in the GMW protocol, which is implemented and further optimized in ABY [70] and ABY2.0 [71]. Such circuits referred to as ‘arithmetic circuits’ in the literature, consist of addition and multiplication ‘gates’, where the values being operated on are defined over a finite field.
Table 3: Studies on cryptography-based ILFPPM.
Paper Year Localization Privacy Method Dataset Metric Adversary Model Attack Model
[44] 2014 kNN Paillier
Simulated
Private
Coms. Overhead
Loc. Error
Fully Malicious Loc. & Data Priv. Attack
[35] 2015 Fuzzy Logic Paillier CRAWDAD Loc. Error Semi-trusted Loc. Priv. Attack
[72] 2016 SVM Paillier Private Coms. Overhead Fully malicious Loc. & Data Priv. Attack
[45] 2018 -
Paillier
Garbled Circuit
- - Semi-trusted
Loc. & Data Priv. Attack
Protocol-based Attack
[43] 2018 -
Paillier
Garbled Circuit
Simulated - Unilateral-malicious
Loc. & Data Priv. Attack
Protocol-based Attack
[73] 2018 kNN
Paillier
Garbled Circuit
Private
JUIndoorLoc
Tampere Uni
Loc. Error
Floor Detection
- Loc. Priv. Attack
[63] 2019 kNN
Paillier
DGK Algorithm
Colombia Uni
Execution Time
Battery Life
[74] 2019 kNN
ABY Framework
Arithmetic Sharing
Boolean Sharing
Yao’s Sharing
Simulated
Private
same as [44]
Coms. Overhead
Run Time
Fully malicious Protocol-based attack
[75] 2019 kNN Trusted Third Party Simulated
Execution Time
Energy Consumption
Fully trusted
Loc. & Data Priv. Attack
Protocol-based Attack
[76] 2019 kNN
Paillier
Spatial Bloom
- Coms. Overhead Semi-trusted Loc. Priv. Attack
[77] 2020 kNN Paillier Private
Coms. Overhead
Loc. Error
Time Cost
Fully malicious Loc. & Data Priv. Attack
[78] 2020 kNN
Expectation-
maximization
Bayes Network
UJIIndoorLoc
Coms. Overhead
Loc. Error
Time Cost
Semi-trusted Loc. Priv. Attack
[64] 2021 kNN
Paillier
One Time Pad
Garbled Circuits
Simulated
Uni of Helsinki
Coms. Overhead
Loc. Error
Time Cost
Semi-trusted Loc. & Data Priv. Attack
[79] 2022 kNN
Arithmetic Sharing
Delta Sharing
Yao’s Sharing
Simulated
Coms. Overhead
Run Time
Semi-trusted -
[80] 2022 MLE Paillier Private
Coms. Overhead
Loc. Error
[81] 2024 kNN
Inner Product
Encryption
Simulated
JUIndoorLoc
Loc. Error
Coms. Overhead
Time Cost
Semi-trusted Data Priv. Attack
5.3.2 Comparative analysis of Cryptographic techniques